1. If Lynn can type a page in p minutes, what piece of the page can she do in 5 minutes?
A. 5/p
B. p - 5
C. p + 5
D. p/5
E. 1- p + 5
Please explain to me how to solve this.

Answer by ptaylor(2198)

Lynn types 1/p pages per min (ck: ((1/p)pages/min)*(p min)=1 page
In 5 minutes, she types (1/pages/min)*5 min=5/p pages
If Lynn can type a page in 2 min, how many pages can she type in 10 min?
First 2 min---1 page
Next 2 min ----1 page
Next 2 min-----1 page
Next 2 min -----1 page
Last 2 min-------1 page
Lynn types at the rate of 1/2 page per min so in 10 min, she types (1/2)*10=5 pages
Hope this helps---ptaylor